OpenStudy (anonymous):

Limit as x approaches infinity of (3x)^(1/3x) step by step help please? :)

OpenStudy (tkhunny):

Introduce a logarithm.

OpenStudy (anonymous):

huh?

OpenStudy (tkhunny):

Rather generally, if \(\lim_{x \rightarrow c}{f(x)} = L, then \lim_{x \rightarrow c}{ln(f(x))} = ln(L)\). Can you use this lovely little hint?

OpenStudy (anonymous):

I need to take the natural log of f(x)?

OpenStudy (tkhunny):

If you want to use my little hint, that will help you get to the natural log of the original limit.

OpenStudy (anonymous):

I'm not understanding.... ;/

OpenStudy (anonymous):

ln((3x^(1/3x))= 3^(1/3x)x^(1/3x)

OpenStudy (tkhunny):

Okay, one clarifying question. Is the exponent \(\dfrac{1}{3x}\) or \(\dfrac{1}{3}x\)?

OpenStudy (tkhunny):

What happened to your logarithm? It was there for a moment.

OpenStudy (anonymous):

huh?

OpenStudy (tkhunny):

Please clarify the original problem statement. \(\lim_{x\rightarrow\infty}(3x)^{1/(3x)}\) or \(\lim_{x\rightarrow\infty}(3x)^{(1/3)x}\)

OpenStudy (anonymous):

^ first one

OpenStudy (tkhunny):

Awesome. We need \(\lim_{x\rightarrow\infty (3x)^{1/(3x)}}\) If we assume that the limit exists, we have \(\lim_{x\rightarrow\infty (3x)^{1/(3x)}} = L\). I just called it"L" because I don't know what it is, yet. Using my little hint, we have: \(\lim_{x\rightarrow\infty} ln\left((3x)^{1/(3x)}\right) = ln(L)\).

OpenStudy (tkhunny):

Make sense so far?

OpenStudy (anonymous):

yeah I understand that.. and ln((3x^((1)/(3x))= 3^((1)/(3x))x^((1)/(3x))

OpenStudy (tkhunny):

No, that's no good. Your logarithm disappeared. \(\ln\left((3x)^{1/(3x)}\right) = \dfrac{1}{3x}ln(3x) = \dfrac{ln(3x)}{3x}\). We're just using logarithm rules.

OpenStudy (anonymous):

ohh... what rule is that specifically? ^

OpenStudy (tkhunny):

Log Rule #3 (They don't really have numbers. :-)) -- \(log\left(a^{b}\right) = b\cdot log(a)\)

OpenStudy (anonymous):

hahah.. then why'd you go from 1/3x*ln(3x) to ln(3x)/3x?

OpenStudy (tkhunny):

That's just multiplication. \(\dfrac{1}{a}\cdot b = \dfrac{b}{a}\).

OpenStudy (anonymous):

hahhaha.. my gosh, i feel stupid! Okay, so now we're at.. \[\lim_{x \rightarrow \infty} \frac{ \ln(3x) }{ 3x }\]

OpenStudy (anonymous):

^ that = 0 But, that's not the answer.. the answer is 1?

OpenStudy (tkhunny):

How did you arive at the conclusion that the limit is zero?

OpenStudy (anonymous):

factoring, using LP rule, and then solving, I got 0?

OpenStudy (tkhunny):

That doesn't quite make sense. Something still going wrong, here, What did you factor?

OpenStudy (anonymous):

I factored out the constant. Maybe I should stop trying to go ahead and just follow lol

OpenStudy (tkhunny):

I was afraid of that. There is no constant to factor out. ln(3x) = ln(3) + ln(x). No factoring available. Okay, enough of that. Let's consider numerator and denominator separately... What happens to the numerator as x increases without bound in the positive direction? What happens to the denominator as x increases without bound in the positive direction?

OpenStudy (anonymous):

what do you mean without bound in the positive direction?

OpenStudy (tkhunny):

These mean about the same thing: 1) x increases without bound in the positive direction 2) \(x\rightarrow\infty\) 3) Some folks might say "x approaches infinity". I just don't like that.

OpenStudy (anonymous):

I'm honestly unsure on how to answer this.. How do I figure it out?

OpenStudy (tkhunny):

If you graphed this, y = 3x, what would it look like?

OpenStudy (anonymous):



OpenStudy (tkhunny):

Close engouh. Now just look at it. As x continues to the right, what will y do?

OpenStudy (anonymous):

goes to infinity, keeps increasing.

OpenStudy (tkhunny):

Perfect. What about y = ln(3x)? x increases and y does what?

OpenStudy (anonymous):

increases to 3 and then goes constant basically?

OpenStudy (tkhunny):

No, that's not right. It also increases without bound, just not as fast, so you might think it levels off. Really, y just keeps getting larger as x increases. Do you believe?

OpenStudy (anonymous):

does it just slowly increase?

OpenStudy (tkhunny):

Using logs base 10 log(1) = 0 log(10) = 1 log(100) = 2 log(1000) = 3 log(10000) = 4 See, it just keeps increasing, albeit WAY slower.

OpenStudy (anonymous):

ohh okay

OpenStudy (tkhunny):

Okay, the point of all this "increase" talk is to consider \(\dfrac{ln(3x)}{3x}\). What we just discovered, was that both numerator and denominator continue to increase. In the notation of Indetermiante Forms, this is type \(\dfrac{\infty}{\infty}\). I'm hoping really hard that you have seen this sort of thing.

OpenStudy (anonymous):

Yeah with LP rule right?

OpenStudy (tkhunny):

Awesome. I like to spell it out, so I can laugh at how close it is to "Hospital". l'hopital's rule woudl suggest we do what with such an indeterminate form?

OpenStudy (anonymous):

It's French ;p we take the derivative of both the denom and num.. (1/x)/(3)

OpenStudy (tkhunny):

Excellent. And what is the limit of that as \(x\rightarrow\infty\)?

OpenStudy (anonymous):

How do I solve for that.. ? I usually put the infinity limits on a calculator, but this gave mme -3E-6...?

OpenStudy (tkhunny):

That's you calculator's way of saying "A really, REALLY small number". What non-negative number can you think of that is less than any other nonnegative number?

OpenStudy (anonymous):

....? I'm confused..

OpenStudy (tkhunny):

The limit of that expression, with increasing x, is zero (0).

OpenStudy (anonymous):

how?

OpenStudy (tkhunny):

How is \(\lim_limits{x\rightarrow\infty}\dfrac{1}{3x} = 0\)? Are you sure you don't want to take that one back? The numerator is constant and the denominator increases.

OpenStudy (anonymous):

no, I get it. But that's not the answer.

OpenStudy (tkhunny):

Yes it is. You're just a little too jumpy. :-) We've been talking so long, it is no surprise that you forgot where we started. We did NOT just say L = 0, we just said ln(L) = 0. Now, go find L.

OpenStudy (anonymous):

No, I've written everything down that we've talked about. and L is the f(x) in the original equation right?

OpenStudy (tkhunny):

Right, and we abandoned that right away. We've been working with ln(f(x)) all along.

OpenStudy (anonymous):

so, now I do ln((3x)^(1/3x))=0?

OpenStudy (tkhunny):

And then get rid of the logarithms. If ln(L) = 0, the \(L = e^{0} = 1\) and we have reached the end of this one.

OpenStudy (anonymous):

How does (3x)^(1/3x)=e^(0)?

OpenStudy (tkhunny):

What we have is a theorem. If lim(f(x)) = L, then lim(ln(f(x)) = ln(L) also If lim(ln(f(x))) = ln(L), then lim(f(x)) = L This is not an algebraic manipulation. This is the application of a theorem.